OpenStudy (anonymous):

Tasha is planning an expansion of a square flower garden in a city park. If each side of the original garden is increased by 7 m, the new total area of the garden will be 144 m². Find the length of each side of the original garden.

OpenStudy (anonymous):

19 m 12 m 5 m √5 m

OpenStudy (anonymous):

x = length of original side of garden so x + 7 is new length and the area = 144 = (x + 7)^2 +- 12 = x + 7 +- 12 - 7 = x x = 5 or - 17 we can not have a length negative so the solution is 5

sam (.sam.):

 For a square, A=Length x Height A=(x+7)(x+7)=144 (x+7)(x+7)=144 (x^(2)+14x+49)=144 x^(2)+14x+49=144 x^(2)+14x-95=0 (x+19)(x-5)=0 x=-19,5 ------------------------- We can't have -19, so the answer is 5